Velocity of projectile when it movesperpendicular to its initial velocity is ucot?

Grade:11

1 Answers

Piyush Kumar Behera
7 years ago
see just make the component of velocity in the horizontal direction constant 
that is vsin\theta=ucos\theta so v=ucot\theta
if you cant understand this ask it here once again.